OpenStudy (anonymous):

graphing form : containing (-2,6) and having slope -3

OpenStudy (anonymous):

does 'graphing form' mean \[y=mx+b\]?

OpenStudy (anonymous):

y=-3x

OpenStudy (anonymous):

yes that's what it means

OpenStudy (anonymous):

here we apply equation (y-y1)=m(x-x1) here, m=-3 and (x1,y1)=(-2,6) therefore, (y-6)=-3(x+2) y=-3x-6+6 so, y =-3x

OpenStudy (anonymous):

this is the graph

OpenStudy (anonymous):

thanks(: how you get the graph on here lol?:p

OpenStudy (anonymous):

i will give you link for that

OpenStudy (anonymous):

here we plot your graph easily

OpenStudy (anonymous):

thanks(: